OpenStudy (anonymous):

the heiht of a cylinder is three times the diameter of the base . The surface area of thre ylinder is 126pi ft^2. What is thr radius of the base?

OpenStudy (anonymous):

Surface area = area of two circles (top and bottom) + circumference of the circle times its height. So, you have: Surface Area = 2*pi*r^2 + 2*pi*r*h Surface Area = 2*pi*r^2 + 2*pi*r*6r Surface Area = 2*pi*r^2 + 12*pi*r^2 = 14*pi*r^2 = 126*pi => r^2 = 126/14 => r^2 = 9 => r = 3